The current price of a stock is $50. The stock is expected to pay a dividend of $1 after 6 months. If the 6-month and 1 yearNEED HELP WITH THSI QUESTION!! WITH EXPLANATIONS
Answer #1

Forward price=(Stock Price-Present value of dividends)*e^(r*forward expiry in years)=(Stock Price-Dividend*e^(-r*time to dividend in years))*e^(r*forward expiry in years)=(50-1*e^(-5%*0.5))*e^(5.5%*1)=51.7965762